The expression of the magnetic field associated with an electromagnetic wave in 'vacuum is given by
B = (100D) ysin(2 x103t +k
Determine the wave number, frequency and the direction of propagation of the wave, " aand the magnitude and direction of the electric field associated with it.
1
Expert's answer
2020-04-27T10:12:26-0400

The wave number is


"k=\frac{\omega }{c}=\frac{2000}{3\cdot10^8}=6.67\cdot10^{-6}m^{-1}"


"f=\frac{2000}{2\pi}=318.3\ Hz""E=Bc=(100)(3\cdot10^8)=3\cdot10^{10}\frac{V}{m}"